OpenStudy (anonymous):

Can someone set up this question for me? I am really struggling/: 5x - 2y + 3z = -23 -x - 2y + 2z = 5 10x + y - 4z = -11 Q1 choose a fav variable A1 Q2 pick a pair of equations, eliminate the fav variable A2 Q3 pick another pair, eliminate that same fav A3 Q4 now you have only 2 variables in 2 equations, like a+b=7 & a-2b=13 A4 Q5 pick a second fav A5 Q6 add equations together and eliminate it A6 Q7 solve for second fav A7 Q8 sub second fav into equation and find last variable A8

OpenStudy (phi):

How about starting with y and the first two equations 5x - 2y + 3z = -23 -x - 2y + 2z = 5 subtract the bottom equation from the top equation can you do that ?

OpenStudy (anonymous):

@phi 4x + 5z = -28 ??

OpenStudy (phi):

the easiest (or at least likely to make mistakes) is to first multiply the 2nd equation by -1 that means multiply both sides and all terms by -1 -1*(-x - 2y + 2z ) = -1*5 distribute the -1. we get \[-1\cdot -x + -1\cdot-2y + -1\cdot2z = -5\\ x+2y -2z = -5 \] now we add, term by term the top equation and this new equation 5x - 2y + 3z = -23 x+ 2y - 2z = -5 what do you get ?

OpenStudy (phi):

can you add 5x+x ?

OpenStudy (anonymous):

@phi 6x

OpenStudy (anonymous):

@phi 6x + z = -28 now what?

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

I dont know what to do now for Q5-Q8 ?? :(

OpenStudy (phi):

ok that looks good. now pick the 2nd and 3rd equations -x - 2y + 2z = 5 10x + y - 4z = -11 and we again want to "get rid of y" multiply the second equation in that pair by 2 (all terms) can you do that ?

OpenStudy (anonymous):

multiply just 10x + y - 4z = -11 by 2 ?

OpenStudy (anonymous):

what do i do with the other? @phi

OpenStudy (triciaal):

You will add it to the new equation you get when you do the multiplication

OpenStudy (triciaal):

so go ahead and do the multiplication and show what you get

OpenStudy (anonymous):

20x + 2y + 8z = -22

OpenStudy (triciaal):

sure he will be back he is going through all the steps with you

OpenStudy (anonymous):

thank you @triciaal is that equation correct?

OpenStudy (triciaal):

not exactly watch your signs

OpenStudy (anonymous):

-8z ?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

so now that i have done that, what do i put into this? Q5 pick a second fav A5 Q6 add equations together and eliminate it A6 Q7 solve for second fav A7 Q8 sub second fav into equation and find last variable A8

OpenStudy (anonymous):

like how do i explain and set that all up into these questions?

OpenStudy (anonymous):

because I have this so far.. does it look correct? 5x - 2y + 3z = -23 -x - 2y + 2z = 5 10x + y - 4z = -11 Q1 choose a fav variable A1 y Q2 pick a pair of equations, eliminate the fav variable A2 5x - 2y + 3z = -23 - -x - 2y + 2z = 5 = 4x + 5z = -28 Q3 pick another pair, eliminate that same fav A3 6x + z = -28 Q4 now you have only 2 variables in 2 equations, like a+b=7 & a-2b=13 A4 4x + 5z = -28 & 6x + z = -28

OpenStudy (phi):

we are doing Q2 we picked the 2nd and 3rd equations -x - 2y + 2z = 5 10x + y - 4z = -11 . You multiplied the 2nd of that pair by 2. you now have the equations -x - 2y + 2z = 5 20x + 2y - 8z = -22 now add those two equations. what do you get?

OpenStudy (phi):

oops , we are doing Q3 Q2 is what we got from the first two equations, up above.

OpenStudy (anonymous):

so Q3 is 6x + z = -28 ??

OpenStudy (phi):

FYI the answer for Q2 Here are the steps 5x - 2y + 3z = -23 -x - 2y + 2z = 5 <-- multiply by -1 and add to the top equation. we get 6x + z = -28

OpenStudy (anonymous):

yeah! that is what I did..

OpenStudy (phi):

for Q3 we picked the 2nd and 3rd equations -x - 2y + 2z = 5 10x + y - 4z = -11 . You multiplied the 2nd of that pair by 2. you now have the equations -x - 2y + 2z = 5 20x + 2y - 8z = -22 now add those two equations. what do you get?

OpenStudy (anonymous):

i got 19x - 6z = -17

OpenStudy (phi):

ok, that is the last part of the answer for Q3

OpenStudy (anonymous):

Q1 choose a fav variable A1 y Q2 pick a pair of equations, eliminate the fav variable A2 5x - 2y + 3z = -23 - -x - 2y + 2z = 5 = 4x + 5z = -28 Q3 pick another pair, eliminate that same fav A3 -x - 2y + 2z = 5 + 20x + 2y - 8z = -22 = 19x - 6z = -17

OpenStudy (phi):

for Q4 we have these equations 6x + z = -28 19x - 6z = -17 for Q5, pick z as the "favorite variable"

OpenStudy (anonymous):

did i put it in those answers correctly so far?

OpenStudy (phi):

looks good.

OpenStudy (phi):

for Q6 6x + z = -28 <--- multiply all terms by 6 19x - 6z = -17

OpenStudy (anonymous):

Q4 now you have only 2 variables in 2 equations, like a+b=7 & a-2b=13 A4 6x + z = -28 & 19x - 6z = -17 Q5 pick a second fav A5 z Q6 add equations together and eliminate it A6 6x + z = -28 + 19x - 6z = -17 = 25x - 5z = -45

OpenStudy (anonymous):

it said add equations together ?

OpenStudy (phi):

unless there is a typo in the original equations, we can't just add them first we multiply the first one by 6 (this will give us a 6z term that will cancel the -6z in the other equation)

OpenStudy (anonymous):

36x + 6z = -168

OpenStudy (anonymous):

multiplied by 6 ^

OpenStudy (phi):

now we can add the equations

OpenStudy (anonymous):

ok..which equations?

OpenStudy (phi):

19x - 6z = -17 36x + 6z = -168

OpenStudy (anonymous):

Q5 pick a second fav A5 z Q6 add equations together and eliminate it A6 36x + 6z = -168 + 19x - 6z = -17 Q7 solve for second fav A7 55x = 185

OpenStudy (anonymous):

is this correct? @phi

OpenStudy (phi):

except -168 + -17 is not 185

OpenStudy (anonymous):

-185 ?

OpenStudy (anonymous):

now what about for Q8 ?

OpenStudy (phi):

yes, so 55x= -185 divide both sides by 55 \[ x = - \frac{185}{55}= - \frac{37}{11} \]

OpenStudy (anonymous):

5x - 2y + 3z = -23 -x - 2y + 2z = 5 10x + y - 4z = -11 Q1 choose a fav variable A1 y Q2 pick a pair of equations, eliminate the fav variable A2 5x - 2y + 3z = -23 - -x - 2y + 2z = 5 = 4x + 5z = -28 Q3 pick another pair, eliminate that same fav A3 -x - 2y + 2z = 5 + 20x + 2y - 8z = -22 = 19x - 6z = -17 Q4 now you have only 2 variables in 2 equations, like a+b=7 & a-2b=13 A4 6x + z = -28 & 19x - 6z = -17 Q5 pick a second fav A5 z Q6 add equations together and eliminate it A6 36x + 6z = -168 + 19x - 6z = -17 x= -37/11 Q7 solve for second fav A7 55x = -185 Q8 sub second fav into equation and find last variable A8

OpenStudy (anonymous):

like that? @phi

OpenStudy (phi):

yes, but when you show the equations, you should put them on separate lines something like this: A6 36x + 6z = -168 + 19x - 6z = -17 ------------------- 55x = -185 x = -185/55 = -37/11

OpenStudy (phi):

for A7, use 19x - 6z = -17 replace x with -37/11

OpenStudy (anonymous):

lol i am confused now ... can u copy and paste that Q&A format in the beginning so i can see?

OpenStudy (phi):

you wrote Q6 add equations together and eliminate it A6 36x + 6z = -168 + 19x - 6z = -17 x= -37/11 it would be clearer if you wrote it as A6 36x + 6z = -168 + 19x - 6z = -17 ------------------- 55x = -185 x = -185/55 = -37/11

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what about Q8 ?

OpenStudy (phi):

for A7, use 19x - 6z = -17 replace x with -37/11

OpenStudy (phi):

are you stuck?

OpenStudy (anonymous):

yes! @phi

OpenStudy (anonymous):

just for Q8 idk what to put

OpenStudy (phi):

first we have to do Q7 Q7 solve for second fav

OpenStudy (phi):

to do that we start with one of the equations from A4 for example, we can use 6x + z = -28 x is -37/11 replace x with that number. we want to find what z is